# On the Capacity of Self-Attention

### v1

@@ Abstract
While self-attention is known to learn relations among tokens, we lack a formal understanding of its capacity: how many distinct relations can a single layer reliably recover for a given budget?
To formalize this, we introduce Relational Graph Recognition (RGR), where the key-query channel represents a graph on $m$ items with $m^{\prime}$ directed edges, and, given a context of items, must recover the neighbors of each item. We measure resources by the total key dimension $D_{K}=h\,d_{k}$. Within this framework, we analytically derive a capacity scaling law and validate it empirically. We show that $D_{K}=\Theta(m^{\prime}\log m^{\prime}/d_{\text{model}})$ is both necessary (information-theoretic lower bound) and sufficient (explicit construction) in a broad class of graphs to recover $m^{\prime}$ relations. This scaling law directly leads to a new, capacity-based rationale for multi-head attention that applies even when each item only attends to a single target. When embeddings are uncompressed ($m=d_{\text{model}}$) and the graph is a permutation, a single head suffices. However, compression ($m>d_{\text{model}}$) forces relations into overlapping subspaces, creating interference that a single large head cannot disentangle. Our analysis shows that allocating a fixed $D_{K}$ across many small heads mitigates this interference, increasing the number of recoverable relations. Controlled single-layer experiments mirror the theory, revealing a sharp performance threshold that matches the predicted capacity scaling and confirms the benefit of distributing $D_{K}$ across multiple heads.
Altogether, these results provide a concrete scaling law for self-attention capacity and a principled design rule for allocating key-query budget across heads.
@@ 1 Introduction
The transformer architecture, and its self-attention mechanism in particular, has revolutionized fields from natural language processing to computer vision [60]. At its core, self-attention computes a similarity-weighted pattern of pairwise relationships among items in a context: queries match keys; the resulting scores route information via values [53, 36, 72, 61, 16, 57]. We ask a basic question: for a fixed self-attention mechanism size, how many target relationships can a single attention layer represent and reliably recover? We call this the layer’s capacity. Capacity is a foundational property for several reasons. (i) It imposes a hard ceiling on relational computation: beyond a threshold, no training procedure or dataset can make a layer uniformly recover all relations, much like rank bounds in linear models. (ii) It complements mechanistic work that isolates specific attention circuits in trained transformers [9, 63, 46, 33], by asking how many independent circuits can coexist. (iii) It provides an actionable resource scaling law, by describing how relationship capacity grows with increasing attention mechanism budget.
@@
We also demonstrate that capacity impacts when increasing the number of heads is useful. Multiple heads are often conceived as a way for a source concept to attend to multiple different targets [60], but looking at self-attention through the lens of capacity demonstrates that multiple heads are beneficial even in the simple case where each concept attends to only a single target. Specifically, when compressed embeddings are used, many relations must be stored in overlapping subspaces; distributing the self-attention budget across many small heads reduces interference and increases the number of relations that can be cleanly separated—consistent with both pruning/specialization studies and expressivity results for attention [64, 45, 10].
@@
One might hope to answer capacity empirically by probing large trained models. In practice, this is ill‑posed. Modern transformers superpose many relationships in shared subspaces; heads are polyfunctional and context‑dependent, so the number of “active” relations is not directly observable. Moreover, attention weights need not align with causal importance [28], and even sophisticated circuit‑tracing pipelines currently miss parts of the QK computation that determine where a head attends [33]. Beyond these methodological issues, superposition makes enumeration intrinsically hard: models can store more features than basis directions, packing multiple concepts into overlapping subspaces [18, 7]. As a result, interpretability work thus far has not revealed how many relationships can be supported by a fixed attention budget.
@@
We therefore introduce a frameworkRelational Graph Recognition (RGR)—and analyze an idealized self‑attention model for solving RGR. The framework allows us to explicitly control both the structure and the number of attention relationships by casting self‑attention as recovering edges of a relational graph among $m$ items, while the model preserves the computational constraints and symmetries of attention. This allows predictions through principled analysis as well as controlled simulations that directly test those predictions. Our abstraction isolates the key–query computation that determines where a head attends, separating it from the OV pathway that determines what is written—a split made explicit in recent mechanistic analyses of attention heads [33]. As a result, our attention budget is defined in terms of the total key dimension $D_{K}=h\,d_{k}$, where $h$ is the number of heads and $d_{k}$ the per‑head key (and query) width.
@@
Let $G=(V,E)$ be a directed graph on $m=\lvert V\rvert$ items with $m^{\prime}=\lvert E\rvert$ edges. A context is an ordered tuple $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$ with $1\leq\ell\leq m$. The Relational Graph Recognition (RGR) mapping takes $(G,\mathcal{C})$ and, for each $v\in\mathcal{C}$, returns its in-context neighbors $N_{G}(v;\mathcal{C})=\{v^{\prime}\in\mathcal{C}:(v,v^{\prime})\in E\}.$ Given a graph $G$, we wish to find a parameterization $\Theta(G)$ such that the mapping defined by $\Theta(G)$ correctly produces $N_{G}(v;\mathcal{C})$ for every context $\mathcal{C}$ and every $v\in\mathcal{C}$.
@@
Fix an input embedding dimension $d_{\text{model}}$. For a graph family $\mathcal{G}_{m,m^{\prime}}=\{\,G:\lvert V\rvert=m,\ \lvert E\rvert=m^{\prime}\},$ we ask for the minimal total key dimension $D_{K}\;=\;h\,d_{k}$ (number of heads times per-head key/query width) such that the self-attention model in Section 3 can realize the RGR mapping for all $G\in\mathcal{G}_{m,m^{\prime}}$ and all contexts $\mathcal{C}$. We refer to this minimal $D_{K}$ as the capacity required by $\mathcal{G}_{m,m^{\prime}}$ at embedding dimension $d_{\text{model}}$.
@@
RGR isolates the key–query channel that determines where attention goes, while preserving the permutation symmetries and parameter sharing of self-attention [^1] We also describe below how positional embeddings can be incorporated into the model.. It lets us dial graph complexity $(m,m^{\prime})$ and the budget $D_{K}$ independently, enabling the information-theoretic bounds, constructive designs, and targeted experiments reported below. Exact mechanics (score computation, aggregation across heads, and omitted components such as softmax and values) are specified in Section 3.
@@
Our analysis yields both fundamental limits and constructive proofs of capability for self-attention as a relational reasoner, and our experiments validate the predictions in an idealized setting that mirrors the theoretical model. The main contributions are:
@@
We cast “where to attend” as Relational Graph Recognition (RGR) and analyze an idealized attention layer that preserves attention symmetries while isolating the key–query computation. The complexity measure is the total key dimension $D_{K}=h\,d_{k}$ (Sec. 3).
@@
We show that recovering $m^{\prime}$-edge graphs on $m$ items with fixed margin requires
$D_{K}=\Omega\!\Big(\tfrac{m^{\prime}}{d_{\text{model}}}\log\tfrac{m^{2}}{m^{\prime}}\Big),$
independent of parameter precision and for any context length $\ell\geq 2$ (Sec. 7). This formalizes the intuition that the key-query matrices need to express sufficient information to describe the underlying relational graph. It also shows that representing more relationships demands greater key–query capacity, and that a smaller embedding dimension requires a larger total key dimension.
@@
We provide explicit attention-based algorithms for RGR within the idealized model (Section 4). We achieve
$D_{K}\;=\;O\!\Big(\tfrac{(m^{\prime}+\Delta)}{d_{\text{model}}}\log m^{\prime}\Big),$
where $\Delta$ is the maximum degree of the graph $G$. Under a mild condition around balanced degrees ($(\Delta/d_{\text{avg}}\leq m/d_{\text{model}})$, this closes the gap between upper and lower bounds for all but very dense graphs, which have a gap factor of $\log m$. These constructions assume random Gaussian unit norm embeddings and also extend to any embedding satisfying a superposition hypothesis style near-orthogonality condition. These constructions also surface the core computational principles of self‑attention and serve as concrete, testable hypotheses about the internal mechanisms transformers might learn.
@@
Our analysis shows that multiple heads are beneficial even when each source has a single correct target (Section 5): when using compressed embeddings $d_{\text{model}}\ll m$, the signals for different relationships are superposed and multiple heads mitigate interference by specializing on disjoint subsets of relationships, thereby allowing the per-head dimension to be small. This provides a principled capacity-centric justification for multi-head attention as a method to reduce noise.
Empirical Validation of Capacity and Head Scaling.
In controlled single-layer experiments that mirror the idealized model, performance exhibits a sharp transition as $D_{K}$ increases (Section 6). The minimal $D_{K}$ needed to reach high accuracy (micro-F1 $\gtrsim 0.99$) grows rapidly with $m$ and shrinks with $d_{\text{model}}$, consistent with the theory. We recover the predicted scaling:
$\boxed{D_{K}^{\star}\;\approx\;C\,\frac{m\log m}{d_{\text{model}}}}$
with a single global threshold at test time. A one‑parameter fit to all data yields $C=1.19$ with $R^{2}=0.944$. Discarding three data points that the theory predicts to perform slightly worse yields $C=0.966$ with $R^{2}=0.992$. We also find a pronounced multi‑head advantage even for permutation graphs: the smallest passing models use several heads while keeping per‑head width small and the optimal head count scales linearly with $\frac{m}{d_{\text{model}}}$, as predicted by the theory. We also see that capacity is largely insensitive to context length over $\ell\!\in\!\{16,32\}$.
Together, these findings yield a quantitative scaling law for capacity as a function of $D_{K}$ and $d_{\text{model}}$, and reveal a principled multi‑head advantage even when each source has only a single correct target—clarifying when and how to allocate key–query budget across heads. The close alignment between our constructive bounds and empirical thresholds provides a concrete, falsifiable foundation for the computational principles that enable self-attention.
@@ 2 Related Work
Closest to our focus are works on memorization capacity [58, 35, 32], including analyses of memorization in attention modules [42]. While aligned in spirit, the problem formulations are different: memorization typically maps each context to a single output token/label, whereas our RGR setting asks for the recovery of in‑context neighbors for every context from a set of possible tokens. Reductions between the two would require memorization handling a combinatorial number of contexts (polynomial in $m$ for fixed $\ell$, and exponential when $\ell$ scales with $m$), and we are not aware of efficient reductions that preserve guarantees in either direction. Accordingly, bounds in one setting do not directly imply bounds in the other. Not surprisingly, capacity results for memorization provide different scaling laws then ours. For example they often include explicit dependence on sequence length $\ell$ [42, 35, 32], whereas in our analysis and experiments, length plays a limited role. Our abstraction isolates the key–query addressing step—“where to attend”—which mechanistic analyses identify as central to head routing [33]. In this sense, RGR complements parameter‑centric memorization settings that emphasize “what to output”: we target the capacity required to select the correct neighbors across contexts.
@@
Beyond memorization, prior theory characterizes what Transformers can compute with sufficient resources—universality/approximation [70], fine‑grained attention‑matrix expressivity [39], and structural bottlenecks such as per‑head low rank and rank collapse without mixing [5, 15]. Language‑theoretic and composition results map limitations at fixed budgets [24, 47]; orthogonally, restricting connectivity rather than dimensions shows universality of $O(\ell)$‑sparse patterns and principled pruning of dense ones [71, 67]; and algorithmic views analyze in‑context procedures [38]. Mechanistic studies find head specialization and prune‑ability [9, 45], while memory‑centric views link attention and FFNs to associative/key–value memories [50, 22]. See Appendix D for more details.

### v3

# A Capacity-Based Rationale for Multi-Head Attention

@@ Abstract
We study the capacity of the self-attention key–query channel: for a fixed budget, how many distinct token–token relations can a single layer reliably encode? We introduce Relational Graph Recognition, where the key–query channel encodes a directed graph and, given a context (a subset of the vertices), must recover the neighbors of each vertex in the context. We measure resources by the total key dimension $D_{K}=h\,d_{k}$. In a tractable multi-head model, we prove matching information-theoretic lower bounds and upper bounds via explicit constructions showing that recovering a graph with $m^{\prime}$ relations in $d_{\text{model}}$-dimensional embeddings requires $D_{K}$ to grow essentially as $m^{\prime}/d_{\text{model}}$ up to logarithmic factors, and we obtain corresponding guarantees for scaled-softmax attention. This analysis yields a new, capacity-based rationale for multi-head attention: even in permutation graphs, where all queries attend to a single target, splitting a fixed $D_{K}$ budget into multiple heads increases capacity by reducing interference from embedding superposition. Controlled experiments mirror the theory, revealing sharp phase transitions at the predicted capacity, and the multi-head advantage persists when adding softmax normalization, value routing, and a full Transformer block trained with frozen GPT-2 embeddings.
@@ 1 Introduction
At the core of the Transformer architecture is self-attention, a mechanism that computes a similarity-weighted pattern of pairwise relationships among items in a context: queries match keys; the resulting scores route information via values Vaswani et al. (2017); Santoro et al. (2017). We ask a basic question: for a fixed self-attention mechanism size, how many target relationships can a single attention layer represent and reliably recover? We call this the layer’s capacity. Capacity is foundational for several reasons. (i) It imposes a hard ceiling on relational computation: beyond a threshold, no training procedure or dataset can make a layer recover all relations, much like rank bounds in linear models. (ii) It complements mechanistic work that isolates specific attention circuits in trained transformers Clark et al. (2019); Vig and Belinkov (2019); Olsson et al. (2022); Kamath et al. (2025), by asking how many independent circuits can coexist and quantifying the interference that limits their coexistence-recently termed the ’Semantic Stability Gap’ Beton and Chana (2026). (iii) It provides an actionable resource scaling law, describing how relationship capacity grows with increasing attention budget.
@@
We also demonstrate that capacity impacts when increasing the number of heads is useful. Multiple heads are often conceived as a way for a source concept to attend to multiple different targets Vaswani et al. (2017), but looking at self-attention through the lens of capacity shows that multiple heads are beneficial even in the simple case where each concept attends to only a single target. Specifically, when compressed embeddings are used, many relations must be stored in overlapping subspaces; distributing the self-attention budget across many small heads reduces interference and increases the number of relations that can be cleanly separated—consistent with both pruning/specialization studies and expressivity results for attention Voita et al. (2019); Michel et al. (2019); Cordonnier et al. (2020b).
@@
One might hope to answer capacity empirically by probing large trained models. In practice, this is ill‑posed. Modern transformers superpose many relationships in shared subspaces; heads are polyfunctional and context‑dependent, so the number of “active” relations is not directly observable. Moreover, attention weights need not align with causal importance Jain and Wallace (2019), and even sophisticated circuit‑tracing pipelines currently miss parts of the QK computation that determine where a head attends Kamath et al. (2025). Beyond these methodological issues, superposition makes enumeration intrinsically hard: models can store more features than basis directions, packing multiple concepts into overlapping subspaces Elhage et al. (2022); Bricken et al. (2023). As a result, interpretability work thus far has not revealed how many relationships can be supported by a fixed attention budget.
@@
We therefore introduce a frameworkRelational Graph Recognition (RGR)—and analyze an idealized self‑attention model for solving RGR. The framework allows us to explicitly control both the structure and the number of attention relationships by casting self‑attention as recovering edges of a relational graph among $m$ items, while the model preserves the computational constraints and symmetries of attention. This allows predictions through principled analysis as well as controlled simulations that directly test those predictions. Our abstraction isolates the key–query computation that determines where a head attends, separating it from the OV pathway that determines what is written—a split made explicit in recent mechanistic analyses of attention heads Kamath et al. (2025). As a result, our attention budget is defined in terms of the total key dimension $D_{K}=h\,d_{k}$, where $h$ is the number of heads and $d_{k}$ the per‑head key (and query) width.
@@
Let $G=(V,E)$ be a directed graph on $m=\lvert V\rvert$ items with $m^{\prime}=\lvert E\rvert$ edges. $G$ encodes the target function to be learned by an attention layer, with vocabulary $V$ and relationships between items $E$. We call an example of that function a context: an ordered subset of distinct vertices $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$ with $1\leq\ell\leq m$. Given a graph $G$, we wish to find a parameterization $\Theta(G)$ of an attention layer such that given any context $\mathcal{C}$, for each $v\in\mathcal{C}$, the attention layer computes its in-context neighbors $N_{G}(v;\mathcal{C})=\{v^{\prime}\in\mathcal{C}:(v,v^{\prime})\in E\}.$ Note that $G$ represents relationships to be learned and not the attention graph; our models always compare all pairs of items in the context.
@@
Fix an input embedding dimension $d_{\text{model}}$. For a graph family $\mathcal{G}_{m,m^{\prime}}=\{\,G:\lvert V\rvert=m,\ \lvert E\rvert=m^{\prime}\},$ we ask for the minimal total key dimension $D_{K}\;=\;h\,d_{k}$ such that the self-attention model in Section 3 can realize the RGR mapping for all $G\in\mathcal{G}_{m,m^{\prime}}$ and all contexts $\mathcal{C}$. We refer to this minimal $D_{K}$ as the capacity required by $\mathcal{G}_{m,m^{\prime}}$ at embedding dimension $d_{\text{model}}$.
@@
RGR isolates the key–query channel that determines where attention goes, while preserving the permutation symmetries and parameter sharing of self-attention [^1] We also describe below how positional embeddings can be incorporated into the model.. It lets us dial graph complexity $(m,m^{\prime})$ and the budget $D_{K}$ independently, enabling the results reported below. Also, given our focus on a rationale for multi-head attention, it allows us to perform analysis and experiments on permutation graphs, where parallel attention to multiple items is not helpful, removing that factor in the advantage of multiple heads. Exact mechanics (score computation and aggregation across heads) are specified in Sections 3 and 8.
@@
Our analysis yields both fundamental limits and constructive proofs of capability for self-attention as a relational reasoner, and our experiments validate these predictions in both the idealized model and more realistic extensions. The main contributions are:
@@
We cast “where to attend” as Relational Graph Recognition (RGR) and analyze two QK variants: the usual scaled-softmax and max-over-heads aggregation, a surrogate for head-wise competition that preserves a key nonlinearity absent from fully linear attention, but is also tractable enough to enable tighter and more extensive analysis (Sec. 3). In both cases, the complexity measure is the total key dimension $D_{K}=h\,d_{k}$.
@@
We prove that recovering graphs with $m^{\prime}$ edges on $m$ items requires a total key dimension of $D_{K}=\Omega\!\left(\tfrac{m^{\prime}}{d_{\text{model}}}\log\tfrac{m^{2}}{m^{\prime}}\right).$ This bound applies to both standard softmax and our max-over-heads variant, quantifying how capacity must grow with the number of relationships, and showing that a smaller model dimension requires a larger total key dimension (Sec. 9).
@@
We provide explicit constructions for RGR. For our max-over-heads model, we achieve $D_{K}=O((\frac{m^{\prime}}{d_{\text{model}}}+\Delta)\log m^{\prime})$, where $\Delta$ is the maximum degree of $G$ (Sec. 4). This is asymptotically optimal for sparse graphs with mild degree imbalance. For softmax attention on permutation graphs, we show $D_{K}=\Theta(\frac{m}{d_{\text{model}}}(\log m)^{2})$ is sufficient (Sec. 8), where the extra $\log$ factor accounts for softmax concentration against distractors. These constructions surface the core computational principles of self‑attention and serve as concrete, testable hypotheses about the internal mechanisms transformers learn.
@@
We show a multi-heads advantage even for permutation graphs (single target per query, and thus no need to attend to multiple targets in parallel - Sec. 5). When $d_{\text{model}}\ll m$, the signals for different relationships are superposed, causing interference which grows with the number of relationships assigned to each head. Thus, more smaller heads outperform fewer larger ones. This noise-reduction mechanism appears under both the max-over-heads and softmax model variants, and provides a principled capacity-centric justification for multi-head attention as a method to reduce noise.
Empirical Validation and Robustness to Model Extensions.
We confirm our theoretical results in controlled single layer experiments:
  • Capacity: We observe sharp phase transitions in performance as $D_{K}$ increases (Sec. 6,7.1,7.3, App. A.1,A.3).
  • Scaling: The minimal budget $D_{K}^{\star}$ follows our predicted scaling for the max-over-heads variant: $D_{K}^{\star}\approx\frac{m^{\prime}\log m}{d_{\text{model}}}$. This trend holds for both permutation graphs and denser regular graphs (Sec. 6).
  • Head Count: The optimal number of heads in our max-over-heads variant scales linearly with $m/d_{\text{model}}$, as predicted theoretically (Sec. 6).
  • Robustness: These trends hold in progressively more realistic model variants: adding softmax (Sec. 7.1), adding an OV/value channel (Sec. 7.2), and even a full single-layer Transformer block trained with frozen GPT-2 embeddings on an induction-style retrieval task only requiring attention to a single target (Sec. 7.3).
Together, these findings give a quantitative and falsifiable picture of how key–query budget enables relational computation in attention: the required total key dimension scales primarily with the number of stored relationships (edges) relative to the embedding dimension, and distributing a fixed budget across multiple heads can be essential even when each query has a single correct target. The alignment between lower bounds, constructive designs, and empirical thresholds persists across denser graphs, softmax normalization, value-routing objectives, and a full Transformer block with realistic frozen embeddings. In addition, our findings are consistent with several widely reported behaviors of attention in LLMs: see Sec. 2.1.
@@ 2 Related Work
Closest to our focus are works on memorization capacity Vardi et al. (2020); Kim et al. (2023); Kajitsuka and Sato (2025), including analyses of memorization in attention modules Mahdavi et al. (2024). While aligned in spirit, the problem formulations are different: memorization typically maps each context to a single output token/label, whereas our RGR setting asks for the recovery of in‑context neighbors for every context from a set of possible tokens. Reductions between the two would require memorization handling a combinatorial number of contexts (polynomial in $m$ for fixed $\ell$, and exponential when $\ell$ scales with $m$), and we are not aware of efficient reductions that preserve guarantees in either direction. Accordingly, bounds in one setting do not directly imply bounds in the other. Not surprisingly, capacity results for memorization provide different scaling laws than ours. Our abstraction isolates the key–query addressing step—“where to attend”—which mechanistic analyses identify as central to head routing Kamath et al. (2025). In this sense, RGR complements parameter‑centric memorization settings that emphasize “what to output”: we target the capacity required to select the correct neighbors across contexts.
@@
Beyond memorization, prior theory characterizes what Transformers can compute with sufficient resources—universality/approximation Yun et al. (2020a), fine‑grained attention‑matrix expressivity Likhosherstov et al. (2021), and structural bottlenecks such as per‑head low rank and rank collapse without mixing Bhojanapalli et al. (2020); Dong et al. (2021). Language‑theoretic and composition results map limitations at fixed budgets Hahn (2020); Peng et al. (2024); orthogonally, restricting connectivity rather than dimensions shows universality of $O(\ell)$‑sparse patterns and principled pruning of dense ones Yun et al. (2020b); Wang et al. (2022); and algorithmic views analyze in‑context procedures Li et al. (2023b). Mechanistic studies find head specialization and prune‑ability Clark et al. (2019); Michel et al. (2019), while memory‑centric views link attention and FFNs to associative/key–value memories Ramsauer et al. (2021); Geva et al. (2021). In recent blog-style reflections on associate memories, Zhong et al. Zhong et al. (2025) demonstrate an exponential gap between linear and softmax attention due to interference. This contrasts with our polynomial gap, highlighting significant differences in problem formulation. They also empirically show that multi-head outperforms single-head attention, and even hypothesize (but do not quantify) this to be due to interference. They do not experiment with single-target scenarios, leaving open the potential role of parallel attention.
2.1 Our findings vs. observations in trained LLMs
Our findings are consistent with several widely reported behaviors of attention in LLMs. Specifically:
Why so many heads?
LLM hyperparameter optimization leads to many heads with modest $d_{k}$ (e.g., GPT-3 175B: 96 heads Brown et al. (2020)). Instead of explaining this through needing to pay attention to dozens of targets simultaneously, and then somehow making use of all the varied information provided by these targets, our theory predicts that increasing head count keeps per head dimension small, thereby decreasing interference.
Similar effects in pre-trained LLMs.
A similar head-count tradeoff and impact of competitive aggregation as we theorize has been observed empirically in GPT-2 training Zhong et al. (2025): when model width is held fixed, collapsing to fewer heads (down to 1) improves a softmax-free linear-kernel attention variant, but degrades standard softmax attention.
Pruning, specialization and redundancy.
Many heads can be pruned with little loss, a small subset is specialized and important, and substantial head redundancy has been found Voita et al. (2019); Michel et al. (2019); Bian et al. (2021); this matches an above-capacity regime where only a fraction of potential relations are heavily used, and is consistent with pressure to increase heads coming from a need to keep per head dimension small, instead of paying parallel attention to many targets. Redundancy of high-value relations across heads can act as noise-reduction, consistent with our theory of interference.
Softmax vs. linear attention.
Linear attention variants often underperform softmax Han et al. (2024); our analysis predicts this because removing competitive gating makes multi-head collapse to an effectively single $D_{K}$-wide linear map, increasing interference.
MQA/GQA as a capacity knob.
Sharing KV heads (MQA) trades quality for speed Shazeer (2019), while restoring multiple KV groups (GQA) recovers quality Ainslie et al. (2023) and is used in deployed models (e.g., Llama 3 Meta AI (2024)); this is consistent with reducing/restoring effective key capacity.
Taken together, these observations span architecture choices (many small heads; MQA/GQA), ablation behavior (pruning/specialization/head merging), representational structure (redundancy), and algorithmic variants (softmax vs. linear). This provides empirical support that the abstractions studied in RGR capture constraints that trained LLMs already operate under.